where to use S proportional to Tsquare and Gallilios odd number ratio in 1D

where to use  S proportional to Tsquare and Gallilios odd number ratio in 1D

Grade:12th pass

1 Answers

moyeen ahmed
5 years ago
we can use the formula s=ut+1/2at^2, when initial velocity of the body is zero and it is moving with an uniform acceleration......................we can use the odd number rule of galileo when a freely falling body is covering equal distances in equal time intervals Let's consider a freely falling body and time intervals of 1 sec. The distance traveled in first interval i.e 1 sec = 1/2 (10) (12) = 5 m The distance traveled in second interval i.e. between 2nd second, = distance traveled in 2 seconds - distance traveled in 1st second = 1/2 (10)(4) - 5 = 15 m Hence we can calculate distance traveled in successive intervals. Now let's try considering ratio of distances in successive intervals, i.e. 15/5 = 3 Similarly if we proceed and find ratio of distance traveled in 3rd second and 2nd second, we'll get 5. This is summarized by Galileo's law of odd numbers, The distances traveled during equal intervals of time, by a body freely falling from rest, stand to one another in the same ratio as the odd numbers beginning with unity.